An electron is accelerated through a potential difference of 2.2 kV and directed into a region between two parallel plates separated by 20 mm with a potential difference of120 V between them. The electron is moving perpendicular to the electric field when it enters the region between the plates. What magnetic field is necessary perpendicular to both the electron path and the electric field so that the electron travels in a straight line?
T
Answer #1

Here,

potential difference , V1 = 2.2 kV

d = 20 mm = 0.020 m

V = 120 V

let the magnetic field needed is B

let the speed of electron after passing trough V1 iw u

Using conservation of energy

0.5 *m * u^2 = V1 * e

0.5 * 9.11 *10^-31 * u^2 = 2.2 *10^3 * 1.602 *10^-19

solving for u

u = 2.78 *10^7 m/s

Now, for

magnetic force = electric force

q * u * B = q * V/d

2.78 *10^7 * B = 120/.020

solving for B

B = 2.16 *10^-4 T

magnetic field is necessary perpendicular to both the electron path and the electric field so that the electron travels in a straight line is 2.16 *10^-4 T